OpenStudy (anonymous):

In a survey, 480 people, or 75%, said they attended a movie at least once a month. How many people were surveyed? A. 640 people B. 360 people C. 555 people D. 405 people

OpenStudy (anonymous):

A. 640

OpenStudy (anonymous):

how do you know though?

OpenStudy (anonymous):

480/.75 = 640

OpenStudy (ddcamp):

If 480 is 75% of the total people surveyed, we can set up the equation: .75 * (Total Surveyed) = 480 Total Surveyed = 480 / .75 = 640

OpenStudy (anonymous):

.75 is equal to 3/4 to wanna find the remaining 1/4 so you would divide

OpenStudy (anonymous):

you*

OpenStudy (anonymous):

= 100 * 480/ 75 = 640 people

OpenStudy (nikato):

Or u can think of it as 75% of x(the total surveyed is 480. As an equation, it'll b 0.75x=480 Divide by sides by 0.75

OpenStudy (anonymous):

oh okay thank you guys for all your help! (:

OpenStudy (anonymous):

Np